OpenStudy (anonymous):

Use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions. x2 = -4x – 4

OpenStudy (anonymous):

im not sure hahaha... sorry ._."

OpenStudy (radar):

The discriminant is the value of\[b ^{2}-4ac\] Do you know what the values of a, b, c are?

OpenStudy (anonymous):

no sorry. never done a problem like this before

OpenStudy (radar):

well a is the coefficient of the x^2 b is the coefficient of x c is the numerical constant your equation is:\[x ^{2}=-4x-4\]

OpenStudy (anonymous):

ok.

OpenStudy (radar):

I am going to rearrange the equation so that appears.\[x ^{2}+4x+4=0\] I added 4x+4 to both sides. This is legal as I did it to both sides of the equal sign. Do you follow that?

OpenStudy (anonymous):

yes

OpenStudy (radar):

a=1 b=4 c=4

OpenStudy (radar):

discriminant is \[1^{2}-(-4)(1)(4)=1+16=17\]

OpenStudy (anonymous):

ok. proceed

OpenStudy (radar):

Now the important partl. If discriminant is positive, there are two real solutions. If ther discriminant is 0, there is one repeated solution. If the discriminant is negative there are no real solutions. Which category did the discriminant for your equation fit ?

OpenStudy (radar):

Hold on a minute, i messed up\[4^{2}-4(1)(4)=16-16=0\]

OpenStudy (anonymous):

ther are two real solutions

OpenStudy (radar):

That was before I corrected the values for the discriminant which now equals 0

OpenStudy (radar):

There is one repeated solution. the equation can be factored (x+2)(x+2)=0 x=-2 x=-2 see one repeated solution.

OpenStudy (radar):

Sorry about the error, but it was noted before closing this thread. good luck in your studies.

OpenStudy (anonymous):

geekgirl1988, what grade are you in? So that we can help accordingly